OpenStudy (anonymous):

which of the following, if any are equavilent to 2 10/11? A. 2+ 10/11 B. 32/11 C. 1 + 21/11

OpenStudy (anonymous):

Your momma so fat....

OpenStudy (anonymous):

I think it's c

OpenStudy (anonymous):

If I understand the question correctly, you want to find which of the fractions equals 2

OpenStudy (anonymous):

Equal to 2 10/11? That part makes no sense..

Directrix (directrix):

@cblrtopas You are in violation of the OpenStudy Code of Conduct "Posting a reply that is not relevant to the topic of discussion within that question.

OpenStudy (texaschic101):

2 10/11 = ((2 x 11 + 10)/11) = 32/11 (a) 2 + 10/11 = 2/1 + 10/11 = 22/11 + 10/11 = 32/11 (correct) (b) 32/11 (correct) (c) 1 + 21/11 = 11/11 + 21/11 = 32/11 (correct) um...they are all equivalent